OpenStudy (anonymous):

HELP! Order pairs. -x+y=-2 2x-3y=4

OpenStudy (anonymous):

do you prefer substitution or elimination ?

OpenStudy (anonymous):

Doesn't matter I'm not quite sure how to solve this..

OpenStudy (compassionate):

Sometimes substitution would be the best choice., elimination would be the best choice.

OpenStudy (anonymous):

lets do substitution is easier... you want to get 1 variable by itself. In this problem, lets take the 1st equation and get y by itself. -x + y = -2 add x to both sides and get y by itself. Can you do that ?

OpenStudy (anonymous):

So, I'll add it also to the -2? I got -1.

OpenStudy (anonymous):

no...let me show you.. -x + x + y = -2 + x the x's on the left side cancel leaving you with y = -2 + x Do you see how I did that ?

OpenStudy (anonymous):

What do I do after that?

OpenStudy (anonymous):

-x + y = -2.....we want the y by itself so we have to move the x to the other side. To do this we do the opposite. What I mean by that is if x is negative. we add x to cancel them out....if x would have been positive, we would subtract x. What you do on one side of the equal sign, you have to do on the other side. so -x + y = -2 (since the x is negative, we add x to both sides because we want the x on the left side to cancel out. -x + x + y = -2 + x. leaving you with y = -2 + x. Is that better ?

OpenStudy (anonymous):

now that we know y = -2 + x...we substitute -2 + x in for y in the other equation....like this : 2x - 3y = 4 2x - 3(-2 + x) = 4 you with me so far ?

OpenStudy (anonymous):

do you understand ?

OpenStudy (anonymous):

Yes, I am! :) Then it'll be 2x+6-3=4 ?

OpenStudy (anonymous):

after multiplying that?

OpenStudy (anonymous):

almost...your forgetting the x's 2x - 3(-2 + x) = 4 2x + 6 - 3x = 4 now we combine like terms -x + 6 = 4 you with me so far ?

OpenStudy (anonymous):

Ahh, I forgot putting the x's lol How would we combine the like terms? Are we going to add 6 & 4 together?

OpenStudy (anonymous):

we are only working on the left side when combining like terms... 2x - 3x = -x....the 6 has no like terms on the left side. so now we have : -x + 6 = 4 now we want to get the 6 on the other side because we want x by itself. We do the opposite. since 6 is positive, we subtract 6 from both sides -x + 6 - 6 = 4 - 6 you understanding this ?

OpenStudy (anonymous):

Ohh, ok I get it know so far!

OpenStudy (anonymous):

I appreciate you helping me out so i'll like to thank you!!

OpenStudy (anonymous):

the 6's on the left side cancel leaving us with -x = -2 We do not want x to be negative so we are going to multiply the equation by -1 to make x positive, but remember what we do on one side we have to do on the other side. (-1)-x = (-1)(-2) understand ?

OpenStudy (anonymous):

Yes, yes I do (:

OpenStudy (anonymous):

by doing that, we get x = 2 you following me ?

OpenStudy (anonymous):

ohhh, okay yeah I am so so far we have x=2 so then we'll sub it onto the other equation? can we do that?

OpenStudy (anonymous):

yes...we can and we will....sub 2 in for x in either of the original equations to find y. Show me what you can do.

OpenStudy (anonymous):

do you need some help ?

OpenStudy (anonymous):

yay!! y=0 ??

OpenStudy (anonymous):

hold on..let me check

OpenStudy (anonymous):

alrighty :D

OpenStudy (anonymous):

2x - 3y = 4 2(2) - 3y = 4 4 - 3y = 4 -3y = 4 - 4 -3y = 0 y = 0/-3 y = 0 yes... your right let us check our answers by subbing in known values into eithr equation and if they come out equal, we have done it right. 2x - 3y = 4 2(2) - 3(0) = 4 4 - 0 = 4 4 = 4 BINGO.....we got it....x = 2 and y = 0 or (2,0) way to go gneyweny :)

OpenStudy (anonymous):

AAAAHHH THANK YOU SO MUCH FOR STICKING AROUND AND HELPING ME OUT! YOU'RE THE BEST! :D

OpenStudy (anonymous):

I don't mind helping when I can.....you did great also :)

OpenStudy (anonymous):

Thank you, again! I really appreciate it! (:

OpenStudy (anonymous):

anytime :)